Sir in this question
Prove that the vectors 3i+2j-2k,-i+3j+4k and 4i-j-6k can form a triangle.
If we find the distance between the vectors and add them then we get zero so it is proved. 
But I cannot understand why do we find the distance between the vectors. Because these already represent vectors which form a triangle and I don't think these are the vertices.
I think I have conceptual misunderstanding.
Pls explain
Thanks
Jai

Grade:12

2 Answers

siddharth gupta
9 years ago
LET a=3i+2j-2k ,b=-i+3j+4k and c=4i-j-6k,you see on adding b and c we obtain a therfore a,b and c form a triangle according to TRIANGLE LAW OF VECTOR ADDITION.
with regards 
SIDDHARTH GUPTA
Ankit Jaiswal
8 years ago
hi
LET a=3i+2j-2k ,b=-i+3j+4k and c=4i-j-6k,you see on adding b and c we obtain a therfore a,b and c form a triangle according to TRIANGLE LAW OF VECTOR ADDITION.
with regards 
SIDDHARTH GUPTA